OpenStudy (adrianna.gongora):

I've been trying this one for hours ;-; can someone please help me I will medal and fan! The shape of a roller coaster is modeled by a polynomial function, R(x). Describe how to find the x-intercepts of R(x) and how to construct a rough graph of R(x) so that the engineer can predict when there will be no change in the direction of the coaster. You may create a sample polynomial to be used in your explanations.

OpenStudy (adrianna.gongora):

Let me explain with this fourth degree of the polynomial. Let's take R(x) = 2x^4 + 3x^3 -11x^2 -9x +15 Now let's find the roots of R(x). By default, let's check with x =1 or x = -1 Let x= 1 R(X) = 2(1)^4 + 3(1)^3 -11(1)^2 - 9(1) + 15 = 2+3-11-9 + 15 = 20 -20 = 0 So R(x) = 0 when x = 1 So x = 1 is one of the root. So (x -1) is a factor. Therefore R(x) = (x -1)(2x^3 + 5x^2 - 6x - 15) To find the other roots, let's use RATIONAL ROOTS test. Let me define the rational root test. The roots will be fractions of the form= ± (factor of the constant term) / (factor of the leading coefficient) Take the constant term. That is the last term. Here the last term is -15. The factors of 15 are 1, 3, 5, and 15 The leading coefficient is 2, so the factors of 2 are 1 and 2

OpenStudy (adrianna.gongora):

what else after this?

OpenStudy (adrianna.gongora):

or is this it?

OpenStudy (adrianna.gongora):

@Nitroster

OpenStudy (adrianna.gongora):

@sleepyjess

OpenStudy (adrianna.gongora):

@sleepyjess

OpenStudy (ac3):

theres a formula to find them i think it's p/q or something like that. I don't remember the formula. You may have it in your notes.

OpenStudy (ac3):

i looked it up it's all of the factors of the last term divided by all of the factors of the leading coefficient

OpenStudy (ac3):

so the easiest way to go about it is to right all of those possibilities and plug them into the equation if you get zero then it's a zero

OpenStudy (adrianna.gongora):

Thank you I have been trying to get help for 3 hours had another post for the same question but no one was helping me so I took it down and made a new one I greatly appreciate your help.

OpenStudy (ac3):

-15: 1,3,5,15 2: 1,2 plus or minus all of these and that should be it.

OpenStudy (ac3):

I have a program that will give me the answer one sec so you can check your work

OpenStudy (adrianna.gongora):

Okay thank you

OpenStudy (ac3):

3/2,1, and it gave me the last 2 as a decimal :(. 1.7321, -1.7321

OpenStudy (ac3):

so one of your fractions is going to equal 1.7321

OpenStudy (ac3):

take this with a grain of salt i'm not even sure anymore whether this is right or not

OpenStudy (ac3):

let me know if you want me to check your work

OpenStudy (adrianna.gongora):

This whole thing is confusing ;-; honestly I don't even know right now.

OpenStudy (ac3):

lol i'll walk you through it

OpenStudy (adrianna.gongora):

Thank you

OpenStudy (ac3):

ok so to start the formula is. (factors of the constant)/(factors of the leading coefficient)

OpenStudy (adrianna.gongora):

were dividing them?

OpenStudy (ac3):

yes

OpenStudy (ac3):

oh and it's plus or minus

OpenStudy (adrianna.gongora):

okay could you set it up for me so I can do it?

OpenStudy (adrianna.gongora):

i'll do the work but what numbers do i plug in lol

OpenStudy (ac3):

so we will have..... \[\pm{1,\frac{ 1 }{ 2 },3,\frac{ 3 }{ 2 },5,\frac{ 5 }{ 2 },15,\frac{ 15 }{ 2 }}\]

OpenStudy (ac3):

you would plug all those numbers in to the equation if you get a zero. then the number you plugged in is a zero

OpenStudy (ac3):

assuming you did everything right up until I showed up everything will work just fine

OpenStudy (adrianna.gongora):

(factors of the constant)/(factors of the leading coefficient) this formula?

OpenStudy (ac3):

yea

OpenStudy (ac3):

factors of the last term divided by the factors of theleading coefficient

OpenStudy (ac3):

your teacher is probably gonna want you to do synthetic or long division at this point though. This is how I do it because it's a lot faster

OpenStudy (adrianna.gongora):

I'll try that real quick. one sec.

OpenStudy (ac3):

my bad i meant to say -5/2

OpenStudy (ac3):

which does work

OpenStudy (ac3):